# Nikhil Kadivar,a‡ Guansheng Li,b‡ Jianlu Zheng,c‡ John M. Higgins,d Ming Dao,c George Em Karniadakis,b and Mengjia Xue

### v1

@@
Nikhil Kadivar,a‡ Guansheng Li,b‡ Jianlu Zheng,c‡ John M. Higgins,d Ming Dao,c George Em Karniadakis,b and Mengjia Xue
@@
footnotetext: a School of Engineering, Brown University, Providence, RI, United Statesfootnotetext: b Division of Applied Mathematics, Brown University, Providence, RI, United Statesfootnotetext: c Department of Materials Science and Engineering, Massachusetts Institute of Technology, Cambridge, MA 02139, United Statesfootnotetext: d Massachusetts General Hospital, Harvard Medical School, Boston, MA 02114, United Statesfootnotetext: e Department of Data Science, Ying Wu College of Computing, New Jersey Institute of Technology, Newark, NJ, United States33footnotetext: These authors contributed equally to this work.11footnotetext: Corresponding authors. E-mail: george_karniadakis@brown.edu; mingdao@mit.edu; mengjia.xu@njit.edu22footnotetext: Electronic Supplementary Information (ESI) available: additional figures and analysis for watershed post-processing.
@@ Introduction
Sickling Kinetics Assay. Centrifuged RBCs were treated with osivelotor (formerly known as GBT440; Pfizer, New York, NY, USA) at 0% (no treatment) and 100% modification levels, based on a 1:1 molar ratio of osivelotor to total hemoglobin. The treated suspensions were incubated at room temperature for 1 h and stored at 4 °C until imaging. Brightfield videos were recorded using a high-resolution CMOS camera (The Imaging Source, Charlotte, NC, USA) mounted on an Olympus X71 inverted microscope (Olympus America, Breinigsville, PA, USA) equipped with a 60× oil-immersion objective lens (NA = 1.25). Recordings were acquired under ambient conditions at 4 frames/s with a resolution of 5472 × 3648 (RGB64). To induce hypoxia, a gas mixture of 2% O2 and 5% CO2 balanced with N2 was introduced into the upper gas channel intersecting the lower cell channel. The RBC suspension (2% hematocrit) was prepared after centrifugation at 1,500 rpm and triple washing with PBS.
@@
To evaluate the predictive accuracy of the proposed framework, we performed controlled experiments under varying cell suspension densities. The corresponding patient sample parameters are summarized in Table 1 Patient I. Using the developed approach, we quantified the time-dependent evolution of the sickle-cell fraction across different suspension densities, as shown in Figure 4. The cell suspension density was defined by the total number of cells detected in each video frame (frame resolution: $5472\times 3648$ pixels). Representative cases with 62 (Video 1), 170 (Video 2), and 417 (Video 3) cells are illustrated in Figure 4(A-C), respectively. For each density level, the left column shows the original image and the middle column shows the segmentation overlay after nnU-Net and watershed-based instance separation; the top and bottom rows correspond to $t=0$ s and $t=120$ s, respectively. In the overlays, green denotes healthy cells, red denotes sickle cells, and yellow highlights overlapping regions that are separated by the watershed post-processing. The rightmost panel illustrates the temporal evolution of sickling dynamics under different suspension densities. The red dashed curves represent manually quantified results and the solid curves denote predictions generated by the combined nnU-Net and watershed framework. The mean absolute error (MAE) in sickle-cell fraction was 0.032, 0.040, and 0.023 for cases 1–3, respectively. While Figure 4 confirms accurate recovery of the sickle-cell fraction over time, Figure S1 (ESI) further demonstrates that watershed-based instance separation improves cell counting accuracy compared with nnU-Net segmentation alone. Overall, the sickling dynamics predicted by the nnU-Net augmented with watershed post-processing exhibit strong agreement with manual quantification, demonstrating the accuracy and robustness of the proposed method across a wide range of cell suspension densities. To avoid overlapping RBCs that can reduce counting accuracy, experiments typically use less densely packed suspensions (e.g., Figure 4B). By allowing the use of denser, overlapping suspensions (e.g., Figure 4C), our method increases the experimental throughput 2.5-fold.
@@
We also tested the proposed framework under different hemoglobin modification levels, as shown in Figure 5, where the time-resolved snapshots of sickling dynamics observed in vitro under varying degrees of hemoglobin modification by osivelotor (0% (Video 4) and 100% (Video 5)). The corresponding patient sample parameters are summarized in Table 1 Patient II. At the onset of the experiment ($t=0$ s), the majority of red blood cells (RBCs) exhibited the characteristic biconcave morphology associated with normoxic conditions. As hypoxia (2% oxygen) progressed over time to $t=120$ s, the RBCs underwent morphological transformations into elongated and crescent-shaped forms, representing various sickling states 18. In contrast, under 100% hemoglobin modification (Figure 5B), the RBC morphology at $t=120$ s remained largely similar to that at $t=0$ s, indicating a substantial inhibition of sickling. Figure 5C compares the predicted sickled fraction from our computational method with that obtained by manual counting. Under 0% hemoglobin modification, the final sickled fractions were approximately 91.0% (prediction) and 91.8% (manual), whereas under 100% modification, the fractions declined markedly to 19.2% (prediction) and 23.5% (manual). The close agreement between the two measurements demonstrates that the proposed framework accurately reproduces experimental observations, confirming its robustness and reliability in quantifying sickling dynamics.
@@
Table 1: Patient information in the experiment.
@@ Acknowledgements
This work was supported by the National Heart, Lung, and Blood Institute of the National Institutes of Health under grant number NIH R01HL154150-06. J.M.H. and M.D. also acknowledge partial support from the NIH under grant number R01HL158102. High performance computing resources were provided by the Center for Computation and Visualization at Brown University.

### v2

# Nikhil Kadivar,a‡ Guansheng Li,b‡ Jianlu Zheng,c‡ Ming Dao,c George Em Karniadakis,b and Mengjia Xud

@@
Nikhil Kadivar,a‡ Guansheng Li,b‡ Jianlu Zheng,c‡ Ming Dao,c George Em Karniadakis,b and Mengjia Xud
@@
footnotetext: a School of Engineering, Brown University, Providence, RI, United Statesfootnotetext: b Division of Applied Mathematics, Brown University, Providence, RI, United Statesfootnotetext: c Department of Materials Science and Engineering, Massachusetts Institute of Technology, Cambridge, MA 02139, United Statesfootnotetext: d Department of Data Science, Ying Wu College of Computing, New Jersey Institute of Technology, Newark, NJ, United States33footnotetext: These authors contributed equally to this work.11footnotetext: Corresponding authors. E-mail: george_karniadakis@brown.edu; mingdao@mit.edu; mengjia.xu@njit.edu22footnotetext: Electronic Supplementary Information (ESI) available: additional figures and analysis for watershed post-processing.
@@ Introduction
Sickling Kinetics Assay. Centrifuged RBCs were treated with osivelotor (formerly known as GBT021601; Pfizer, New York, NY, USA) at 0% (no treatment) and 100% modification levels, based on a 1:1 molar ratio of osivelotor to total hemoglobin. The treated suspensions were incubated at room temperature for 1 h and stored at 4 °C until imaging. Brightfield videos were recorded using a high-resolution CMOS camera (The Imaging Source, Charlotte, NC, USA) mounted on an Olympus X71 inverted microscope (Olympus America, Breinigsville, PA, USA) equipped with a 60× oil-immersion objective lens (NA = 1.25). Recordings were acquired under ambient conditions at 4 frames/s with a resolution of 5472 × 3648 (RGB64). To induce hypoxia, a gas mixture of 2% O2 and 5% CO2 balanced with N2 was introduced into the upper gas channel intersecting the lower cell channel. The RBC suspension (2% hematocrit) was prepared after centrifugation at 1,500 rpm and triple washing with PBS.
@@
To evaluate the predictive accuracy of the proposed framework, we performed controlled experiments under varying cell suspension densities. The corresponding patient sample parameters are summarized in Table 1 Sample I. Using the developed approach, we quantified the time-dependent evolution of the sickle-cell fraction across different suspension densities, as shown in Figure 4. The cell suspension density was defined by the total number of cells detected in each video frame (frame resolution: $5472\times 3648$ pixels). Representative cases with 62 (Video 1), 170 (Video 2), and 417 (Video 3) cells are illustrated in Figure 4(A-C), respectively. For each density level, the left column shows the original image and the middle column shows the segmentation overlay after nnU-Net and watershed-based instance separation; the top and bottom rows correspond to $t=0$ s and $t=120$ s, respectively. In the overlays, green denotes healthy cells, red denotes sickle cells, and yellow highlights overlapping regions that are separated by the watershed post-processing. The rightmost panel illustrates the temporal evolution of sickling dynamics under different suspension densities. The red dashed curves represent manually quantified results and the solid curves denote predictions generated by the combined nnU-Net and watershed framework. The mean absolute error (MAE) in sickle-cell fraction was 0.032, 0.040, and 0.023 for cases 1–3, respectively. While Figure 4 confirms accurate recovery of the sickle-cell fraction over time, Figure S1 (ESI) further demonstrates that watershed-based instance separation improves cell counting accuracy compared with nnU-Net segmentation alone. Overall, the sickling dynamics predicted by the nnU-Net augmented with watershed post-processing exhibit strong agreement with manual quantification, demonstrating the accuracy and robustness of the proposed method across a wide range of cell suspension densities. To avoid overlapping RBCs that can reduce counting accuracy, experiments typically use less densely packed suspensions (e.g., Figure 4B). By allowing the use of denser, overlapping suspensions (e.g., Figure 4C), our method increases the experimental throughput 2.5-fold.
@@
We also tested the proposed framework under different hemoglobin modification levels, as shown in Figure 5, where the time-resolved snapshots of sickling dynamics observed in vitro under varying degrees of hemoglobin modification by osivelotor (0% (Video 4) and 100% (Video 5)). The corresponding patient sample parameters are summarized in Table 1 Sample II. At the onset of the experiment ($t=0$ s), the majority of red blood cells (RBCs) exhibited the characteristic biconcave morphology associated with normoxic conditions. As hypoxia (2% oxygen) progressed over time to $t=120$ s, the RBCs underwent morphological transformations into elongated and crescent-shaped forms, representing various sickling states 18. In contrast, under 100% hemoglobin modification (Figure 5B), the RBC morphology at $t=120$ s remained largely similar to that at $t=0$ s, indicating a substantial inhibition of sickling. Figure 5C compares the predicted sickled fraction from our computational method with that obtained by manual counting. Under 0% hemoglobin modification, the final sickled fractions were approximately 91.0% (prediction) and 91.8% (manual), whereas under 100% modification, the fractions declined markedly to 19.2% (prediction) and 23.5% (manual). The close agreement between the two measurements demonstrates that the proposed framework accurately reproduces experimental observations, confirming its robustness and reliability in quantifying sickling dynamics.
@@
Table 1: Clinical information about samples in the experiment.
@@ Acknowledgements
This work was supported by the National Heart, Lung, and Blood Institute of the National Institutes of Health under grant number NIH R01HL154150-06. J.M.H. and M.D. also acknowledge partial support from the NIH under grant number R01HL158102. We thank Dr. John M. Higgins for providing the sickle cell samples. High-performance computing resources were provided by the Center for Computation and Visualization at Brown University. Osivelotor used in this study was provided by Pfizer Inc. through a Pure Compound Grant. Global Blood Therapeutics, the original developer of osivelotor, was wholly acquired by Pfizer Inc. in 2023. We gratefully acknowledge Pfizer Inc. for providing the compound and supporting this research.